# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v2

@@ Abstract
Recent generative models based on score matching and flow matching have significantly advanced generation tasks, but their potential in discriminative tasks remains underexplored. Previous approaches, such as generative classifiers, have not fully leveraged the capabilities of these models for discriminative tasks due to their intricate designs. We propose Pretrained Reversible Generation (PRG), which extracts unsupervised representations by reversing the generative process of a pretrained continuous generation model. PRG effectively reuses unsupervised generative models, leveraging their high capacity to serve as robust and generalizable feature extractors for downstream tasks. This framework enables the flexible selection of feature hierarchies tailored to specific downstream tasks. Our method consistently outperforms prior approaches across multiple benchmarks, achieving state-of-the-art performance among generative model based methods, including 78% top-1 accuracy on ImageNet at a resolution of 64×64. Extensive ablation studies, including out-of-distribution evaluations, further validate the effectiveness of our approach.
@@ 1 Introduction
Generative models formulated as continuous-time stochastic differential equations, including diffusion and flow models [36, 38, 50, 2], have shown remarkable proficiency in multi-modal content generation tasks [52, 27, 44], as well as in scientific modeling [1]. By effectively learning high-dimensional distributions, these models can generate new, high-quality samples that resemble the original data.
@@
While generative tasks require models to reconstruct the entire data distribution, discriminative tasks (e.g., image classification) rely on learning representations where some parts capture the essential underlying data structure to produce discrete or continuous predictions. Previous works [17, 23] have explored various approaches to enhance representation learning, with growing focus on unsupervised visual representation learning [12, 10]. These methods construct generic proxy tasks to efficiently extract and utilize data for pre-training, yielding robust representations for diverse downstream tasks [5, 51, 13, 59, 16, 10, 65].
@@
Due to the extraordinary performance of diffusion models in generative tasks, recent studies have explored their potential for discriminative tasks [14, 34]. These studies often rely on extracting or combining representations from the internal layers of pretrained diffusion models, involving intricate and poorly generalizable designs, yet usually lacking a clear explanation. For example, [65] selects a specific layer in a UNet-based diffusion model as the feature representation. Moreover, the performance of these methods exhibits significant room for improvement compared to current state-of-the-art techniques.
@@
Inspired by the reversible property [11, 56] of continuous-time stochastic flow models, we introduce Pretrained Reversible Generation as Unsupervised Visual Representation Learning (PRG). Our approach applies generative models to discriminative tasks by reversing the generative process into a meaningful inference process. Specifically, we pretrain a reversible generative model via flow matching to maximize the lower bound of mutual information between the original image and its optimal representation, providing a strong initialization. We then reverse the generative process and leverage it for downstream training to improve task performance. This model-agnostic method does not require access to any internal features of a pretrained generative model, eliminating the need for fixed network modules to capture data representations. As a pretrain-finetune paradigm, it fully leverages the powerful capabilities of various pretrained generative models (e.g., diffusion models, flow models). During the reversed process, it utilizes features from different hierarchical levels, enabling efficient adaptation to downstream tasks. Through theoretical analysis Sec. 3.2 and verification experiments Sec. 4.3, we explain why reversing a pretrained generative process serves as an effective feature extractor for downstream tasks. Our main experiments further demonstrate that this method achieves competitive performance in image classification and OOD detection while effectively adapting community-pretrained text-to-image models for downstream training.
@@ 2 Related Work
Generative classifiers is first proposed in [34, 9, 14, 8, 73, 49] model the conditional distribution $p(x|y)$ and use Bayes’ rule to compute the probability $p(y|x)=\frac{p(x|y)p(y)}{p(x)}$. These classifiers are typically trained by maximizing the conditional logarithmic likelihood: $\max_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}\log p(x|y)$.
When using conditional diffusion models, $p(x|y)$ can be computed with the instantaneous change-of-variable formula [11, 22] or approximated by ELBO [55, 26]. Duvenaud et al. [18] advocate an energy-based model to improve calibration, robustness, and out-of-distribution detection. Yang et al. [67] propose HybViT, which combines diffusion models and discriminative classifiers to predict $p(y|x)$. Zimmermann et al. [73] introduce score-based generative classifiers but find them vulnerable to adversarial attacks. Clark and Jaini [14] and Li et al. [34] explore zero-shot generative classifiers using pretrained models like Stable Diffusion [52] to denoise images based on text labels.
In contrast, our approach fine-tunes a classifier on top of the generative model, making it a discriminative classifier that directly estimates $p(y|x)$. Discriminative classifiers are often preferred for classification tasks [60, 47]. Fetaya et al. [19] suggest a trade-off between likelihood-based modeling and classification accuracy, implying that improving generation may hinder classification for generative classifier. However, our two-stage training scheme shows that enhancing the generative model in the pretraining stage can improve discriminative performance during the fine-tuning stage. This suggests a complementary relationship between likelihood-based training and classification accuracy, indicating that discriminative models can benefit from generative pretraining. Thus, we demonstrate a viable approach to leveraging generative models for discriminative tasks.
@@
Generative models have been widely explored for visual representation learning [29, 59, 21]. Denoising autoencoders (DAEs) [62, 63, 6, 20] learn robust representations by reconstructing corrupted input data, which can be interpreted through a manifold learning perspective [61]. Our method extends Vincent et al. [63] by leveraging reversible flow-based generative models like diffusion models.
Diffusion models [55, 26, 57] have shown promise in generation, yet their potential for feature extraction is underexplored. Chen et al. [10] pretrain a Transformer for autoregressive pixel prediction, demonstrating competitive results in representation learning. Their two-stage approach (pre-training a generative model followed by full fine-tuning) closely resembles ours. Additionally, their downsampling of images to low resolution mimics the effect of adding noise, analogous to denoising. However, their model and autoregressive generation scheme are unsuitable for reverse generation and lack efficiency. He et al. [24] pretrain a masked autoencoder, effectively performing denoising by reconstructing incomplete data. Baranchuk et al. [4] show diffusion models improve semantic segmentation, particularly with limited labeled data. Sinha et al. [54] propose Diffusion-Decoding with contrastive learning to enhance representation quality, while Mittal et al. [45] introduce a Diffusion-based Representation Learning (DRL) framework, leveraging denoising score matching.
@@ 3.1 Pretrained Reversible Generation Framework
A diffusion model reverses a forward diffusion process described by $p(x_{t}\mid x_{0})\sim\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}I)$, where $\alpha_{t}$ is a scaling factor and $\sigma_{t}^{2}$ denotes the variance. We adopt the Generalized VP-SDE, which uses a triangular function as coefficients to define the diffusion path, referred to as PRG-GVP with $\alpha_{t}=\cos\left(\frac{1}{2}\pi t\right)$ and $\sigma_{t}=\sin\left(\frac{1}{2}\pi t\right)$. Due to the Fokker-Planck equation, an equivalent ODE can be derived that shares the same marginal distribution, known as the Probability Flow ODE [56]:
@@
A flow model deterministically transforms $x_{0}$ to $x_{1}$ via a parameterized velocity function $v(x_{t})$. Inspired by [58], we adopt the flow path formulation, referred to as PRG-ICFM, with $p(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t}|tx_{1}+(1-t)x_{0})$, and velocity function: $v(x_{t}|x_{0},x_{1})=x_{1}-x_{0}$.
We also employ the Optimal Transport Conditional Flow Matching (OTCFM) model [58]. Specifically, we jointly sample $(x_{0},x_{1})$ from the optimal transport plan $\pi$ to construct the 2-Wasserstein target distribution. The corresponding velocity field $v(x_{t}\mid x_{0},x_{1})$ is then used to approximate dynamic optimal transport, referred to as PRG-OTCFM. In practice, we parameterize the velocity model $v(x_{t})$ as $v_{\theta}$ using a network to model the above variants of PRG.
@@ 3.2 Pretrained Reversible Generation Mechanism
In the traditional autoencoder paradigm, representation learning aims to learn a network that maps the input data $X$ into a useful representation $Z$ for downstream tasks. We parameterize the encoder $p_{\theta}(z|x)$ with parameters $\theta$. A good representation should retain sufficient information about the input $X$. In information-theoretic terms [35, 63], this corresponds to maximizing the mutual information $\mathcal{I}(X,Z)$ between the input random variable $X$ and its representation $Z$. Following the reasoning of Vincent et al. [63], we show that pre-training diffusion or flow models maximizes the mutual information. Specifically, we aim to find the optimal $\theta^{*}$ that maximizes $\mathcal{I}(X,Z)$:
@@
In continuous-time stochastic flow generative models, the inference and generation process are reversible via a same model, which means the decoder parameter $\theta^{\prime}$ and encoder parameter $\theta$ are shared parameters. However, we still use these two notation in the later part of this paper for distinguishing the inference and generation process through ODE solver with $\theta$ and $\theta^{\prime}$ respectively. For preventing any confusion, we denote $x_{0}$ being data variable $x$ amd $x_{1}$ being the inferred latent variable $z$. In practice, training generative models using score matching or flow matching objectives does not directly maximize the likelihood of reconstructing $x_{0}$ from $x_{1}$. However, as shown in previous works [57, 41, 72], these methods provide a lower bound on the likelihood. Specifically, they establish that:
@@
Generative features may not be optimal for discriminative tasks, as reconstruction requires more detail than comprehension. Hence, fine-tuning is necessary to adapt the model for downstream tasks. A key question is whether the flow model should also be fine-tuned alongside the classifier. Due to the flow model’s large capacity, reusing the pretrained model seems preferable. Initially, we explored freezing the model while fine-tuning only the classifier. However, this approach underperformed, indicating that fine-tuning both the flow model and the classifier is necessary. Otherwise, a large and complex classifier would be required, reducing efficiency. To fine-tune the entire model, we leverage neural ODEs [11] for efficient gradient computation, optimizing both $\theta$ and $\phi$ via backpropagation. We observed that distinct discriminative tasks correspond to different parts of generative features Appendix D. Thus, fine-tuning can selectively prune or enhance relevant features as needed. In addition, we assess the robustness and generalization of these features against out-of-distribution samples through further experiments.
@@ 3.3 Advantages
Our approach is agnostic to the choice of network architecture for the generative model: it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models [56]. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [65] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
@@
PRG leverages the infinite-layer structure of continuous-time stochastic flow models [11], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like SD3 [52] and DALL·E 3 [7]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering excellent performance with enhanced capacity.
@@
Features extracted by reversing the generative process remain stable along the reverse trajectory, allowing feature use from any point $t\in[0,1]$ for downstream tasks, not just $x_{1}$. These features are robust to various discretization schemes (e.g., Euler, Runge-Kutta) and time steps ($10^{3}$ to $10^{1}$). Our method generalizes well across datasets, enabling fine-tuning of pretrained models for new tasks. Notably, it applies to all diffusion and flow models and has been successfully transferred to community-developed text-to-image models in Sec. 4.4.3.
In summary, our approach bridges the gap between generative and discriminative models, offering a simple yet effective way to leverage pretrained generative models for unsupervised representation learning, with the ability to fine-tune for downstream supervised tasks. A more detailed analysis of why PRG is effective can be found in Appendix A.
@@ 4.1 Settings
For generative model selection, we adopt three diffusion and flow-based models: GVP, ICFM, and OTCFM. For pre-training, we followed the protocol from [48], using the Adam optimizer [30] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [39], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. We followed DDAE [65]’s setting (64 resolution, simple data augmentation, no mixup [71], no cutmix [69]) for fair comparison, yet our method achieves SOTA diffusion classifier performance. Further details, including hyperparameters and efficiency statistics for training and inference, are provided in Secs. B.1, B.2 and B.3 .
@@
We select state-of-the-art (SOTA) representation models as baselines to evaluate PRG on CIFAR-10 [32], Tiny-ImageNet [33], and ImageNet [53]. Baselines include classical discriminative models like WideResNet-$28$-$10$[70] and ResNeXt-$29$-$16\times 64$d[66], as well as diffusion-based generative models such as GLOW [19], Energy Model [18], SBGC [73], HybViT [67], and DDAE [65]. For out-of-distribution (OOD) evaluation, we compare PRG with PGD [43], PLAT [64], AugMix [25], AutoAug [15], SBGC, and PDE+ [68] on CIFAR-10-C and Tiny-ImageNet-C. All results are from their original papers.

### v3

@@ Abstract
Recent generative models based on score matching and flow matching have significantly advanced generation tasks, but their potential in discriminative tasks remains underexplored. Previous approaches, such as generative classifiers, have not fully leveraged the capabilities of these models for discriminative tasks due to their intricate designs. We propose Pretrained Reversible Generation (PRG), which extracts unsupervised representations by reversing the generative process of a pretrained continuous generation model. PRG effectively reuses unsupervised generative models, leveraging their high capacity to serve as robust and generalizable feature extractors for downstream tasks. This framework enables the flexible selection of feature hierarchies tailored to specific downstream tasks. Our method consistently outperforms prior approaches across multiple benchmarks, achieving state-of-the-art performance among generative model based methods, including 78% top-1 accuracy on ImageNet at a resolution of 64×64. Extensive ablation studies, including out-of-distribution evaluations, further validate the effectiveness of our approach. Code is available at https://github.com/opendilab/PRG.
@@ 1 Introduction
“What I cannot create, I do not understand.”
— Richard Feynman
@@
Generative models formulated as continuous-time stochastic differential equations, including diffusion and flow models [lipman2023flow, liu2022flow, pmlr-v202-pooladian23a, Albergo2023], have shown remarkable proficiency in multi-modal content generation tasks [rombach2022high, Ho2022video, mittal2021symbolic], as well as in scientific modeling [abramson2024accurate]. By effectively learning high-dimensional distributions, these models can generate new, high-quality samples that resemble the original data.
@@
While generative tasks require models to reconstruct the entire data distribution, discriminative tasks (e.g., image classification) rely on learning representations where some parts capture the essential underlying data structure to produce discrete or continuous predictions. Previous works [dosovitskiy2020image, he2016deep] have explored various approaches to enhance representation learning, with growing focus on unsupervised visual representation learning [chen2020simple, chen2020generative]. These methods construct generic proxy tasks to efficiently extract and utilize data for pre-training, yielding robust representations for diverse downstream tasks [bengio2013representation, radford2015unsupervised, chen2016infogan, van2017neural, donahue2019large, chen2020generative, xiang2023denoising].
@@
Due to the extraordinary performance of diffusion models in generative tasks, recent studies have explored their potential for discriminative tasks [clark2024text, li2023your]. These studies often rely on extracting or combining representations from the internal layers of pretrained diffusion models, involving intricate and poorly generalizable designs, yet usually lacking a clear explanation. For example, [xiang2023denoising] selects a specific layer in a UNet-based diffusion model as the feature representation. Moreover, the performance of these methods exhibits significant room for improvement compared to current state-of-the-art techniques.
@@
Inspired by the reversible property [Chen2018, song2020score] of continuous-time stochastic flow models, we introduce Pretrained Reversible Generation as Unsupervised Visual Representation Learning (PRG). Our approach applies generative models to discriminative tasks by reversing the generative process into a meaningful inference process. Specifically, we pretrain a reversible generative model via flow matching to maximize the lower bound of mutual information between the original image and its optimal representation, providing a strong initialization. We then reverse the generative process and leverage it for downstream training to improve task performance. This model-agnostic method does not require access to any internal features of a pretrained generative model, eliminating the need for fixed network modules to capture data representations. As a pretrain-finetune paradigm, it fully leverages the powerful capabilities of various pretrained generative models (e.g., diffusion models, flow models). During the reversed process, it utilizes features from different hierarchical levels, enabling efficient adaptation to downstream tasks. Through theoretical analysis Sec. 3.2 and verification experiments Sec. 4.3, we explain why reversing a pretrained generative process serves as an effective feature extractor for downstream tasks. Our main experiments further demonstrate that this method achieves competitive performance in image classification and OOD detection while effectively adapting community-pretrained text-to-image models for downstream training.
@@ 2 Related Work
Generative classifiers, first proposed in [li2023your, chen2023robust, clark2024text, chendiffusion, zimmermann2021score, pmlr-v162-nie22a], model the conditional distribution $p(x\mid y)$ and use Bayes’ rule to compute $p(y\mid x)=\frac{p(x\mid y),p(y)}{p(x)}$. These classifiers are typically trained by maximizing the conditional logarithmic likelihood: $\max_{\theta}\mathbb{E}_{(x,y)\sim p(x,y)}\log p(x|y)$.
When using conditional diffusion models, $p(x|y)$ can be computed with the instantaneous change-of-variable formula [Chen2018, Grathwohl2018] or approximated by ELBO [sohl2015deep, ho2020denoising]. duvenaud2020your advocate an energy-based model to improve calibration, robustness, and out-of-distribution detection. yang2022your propose HybViT, which combines diffusion models and discriminative classifiers to predict $p(y|x)$. zimmermann2021score introduce score-based generative classifiers but find them vulnerable to adversarial attacks. clark2024text and li2023your explore zero-shot generative classifiers using pretrained models like Stable Diffusion [rombach2022high] to denoise images based on text labels.
In contrast, our approach fine-tunes a classifier on top of the generative model, making it a discriminative classifier that directly estimates $p(y|x)$. Discriminative classifiers are often preferred for classification tasks [Vapnik1995, AndrewNIPS2001]. fetaya2019understanding suggest a trade-off between likelihood-based modeling and classification accuracy, implying that improving generation may hinder classification for generative classifier. However, our two-stage training scheme shows that enhancing the generative model in the pretraining stage can improve discriminative performance during the fine-tuning stage. This suggests a complementary relationship between likelihood-based training and classification accuracy, indicating that discriminative models can benefit from generative pretraining. Thus, we demonstrate a viable approach to leveraging generative models for discriminative tasks.
@@
Generative models have been widely explored for visual representation learning [kingma2013auto, van2017neural, goodfellow2020generative]. Denoising autoencoders (DAEs) [vincent2008extracting, vincent2010stacked, bengio2013generalized, geras2014scheduled] learn robust representations by reconstructing corrupted input data, which can be interpreted through a manifold learning perspective [vincent2011]. Our method extends vincent2010stacked by leveraging reversible flow-based generative models like diffusion models.
Diffusion models [sohl2015deep, ho2020denoising, song2021maximum] have shown promise in generation, yet their potential for feature extraction is underexplored. chen2020generative pretrain a Transformer for autoregressive pixel prediction, demonstrating competitive results in representation learning. Their two-stage approach (pre-training a generative model followed by full fine-tuning) closely resembles ours. Additionally, their downsampling of images to low resolution mimics the effect of adding noise, analogous to denoising. However, their model and autoregressive generation scheme are unsuitable for reverse generation and lack efficiency. he2022masked pretrain a masked autoencoder, effectively performing denoising by reconstructing incomplete data. baranchuk2021label show diffusion models improve semantic segmentation, particularly with limited labeled data. sinha2021dc propose Diffusion-Decoding with contrastive learning to enhance representation quality, while mittal2023diffusion introduce a Diffusion-based Representation Learning (DRL) framework, leveraging denoising score matching.
@@ 3.1 Pretrained Reversible Generation Framework
A diffusion model reverses a forward diffusion process described by $p(x_{t}\mid x_{0})\sim\mathcal{N}(x_{t}\mid\alpha_{t}x_{0},\sigma_{t}^{2}I)$, where $\alpha_{t}$ is a scaling factor and $\sigma_{t}^{2}$ denotes the variance. We adopt the Generalized VP-SDE, which uses a triangular function as coefficients to define the diffusion path, referred to as PRG-GVP with $\alpha_{t}=\cos\left(\frac{1}{2}\pi t\right)$ and $\sigma_{t}=\sin\left(\frac{1}{2}\pi t\right)$. Due to the Fokker-Planck equation, an equivalent ODE can be derived that shares the same marginal distribution, known as the Probability Flow ODE [song2020score]:
@@
A flow model deterministically transforms $x_{0}$ to $x_{1}$ via a parameterized velocity function $v(x_{t})$. Inspired by [tong2024improving], we adopt the flow path formulation, referred to as PRG-ICFM, with $p(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t}|tx_{1}+(1-t)x_{0})$, and velocity function: $v(x_{t}|x_{0},x_{1})=x_{1}-x_{0}$.
We also employ the Optimal Transport Conditional Flow Matching (OTCFM) model [tong2024improving]. Specifically, we jointly sample $(x_{0},x_{1})$ from the optimal transport plan $\pi$ to construct the 2-Wasserstein target distribution. The corresponding velocity field $v(x_{t}\mid x_{0},x_{1})$ is then used to approximate dynamic optimal transport, referred to as PRG-OTCFM. In practice, we parameterize the velocity model $v(x_{t})$ as $v_{\theta}$ using a network to model the above variants of PRG.
@@ 3.2 Pretrained Reversible Generation Mechanism
In the traditional autoencoder paradigm, representation learning aims to learn a network that maps the input data $X$ into a useful representation $Z$ for downstream tasks. We parameterize the encoder $p_{\theta}(z|x)$ with parameters $\theta$. A good representation should retain sufficient information about the input $X$. In information-theoretic terms [linsker1988application, vincent2010stacked], this corresponds to maximizing the mutual information $\mathcal{I}(X,Z)$ between the input random variable $X$ and its representation $Z$. Following the reasoning of vincent2010stacked, we show that pre-training diffusion or flow models maximizes the mutual information. Specifically, we aim to find the optimal $\theta^{*}$ that maximizes $\mathcal{I}(X,Z)$:
@@
In continuous-time stochastic flow generative models, the inference and generation process are reversible via a same model, which means the decoder parameter $\theta^{\prime}$ and encoder parameter $\theta$ are shared parameters. However, we still use these two notation in the later part of this paper for distinguishing the inference and generation process through ODE solver with $\theta$ and $\theta^{\prime}$ respectively. For preventing any confusion, we denote $x_{0}$ being data variable $x$ amd $x_{1}$ being the inferred latent variable $z$. In practice, training generative models using score matching or flow matching objectives does not directly maximize the likelihood of reconstructing $x_{0}$ from $x_{1}$. However, as shown in previous works [song2021maximum, lu2022maximum, zheng2023improved], these methods provide a lower bound on the likelihood. Specifically, they establish that:
@@
Generative features may not be optimal for discriminative tasks, as reconstruction requires more detail than comprehension. Hence, fine-tuning is necessary to adapt the model for downstream tasks. A key question is whether the flow model should also be fine-tuned alongside the classifier. Due to the flow model’s large capacity, reusing the pretrained model seems preferable. Initially, we explored freezing the model while fine-tuning only the classifier. However, this approach underperformed, indicating that fine-tuning both the flow model and the classifier is necessary. Otherwise, a large and complex classifier would be required, reducing efficiency. To fine-tune the entire model, we leverage neural ODEs [Chen2018] for efficient gradient computation, optimizing both $\theta$ and $\phi$ via backpropagation. We observed that distinct discriminative tasks correspond to different parts of generative features Appendix D. Thus, fine-tuning can selectively prune or enhance relevant features as needed. In addition, we assess the robustness and generalization of these features against out-of-distribution samples through further experiments.
@@ 3.3 Advantages
Our approach is agnostic to the choice of network architecture for the generative model: it can be a U-Net, Transformer, or any other model. The latent variable $Z$ remains stationary with respect to the data $X$ when using an ODE solver due to the static nature of continuous-time stochastic flow models [song2020score]. This allows different architectures to encode the data while yielding the same latent variable $Z$. In contrast, methods like DDAE [xiang2023denoising] are model-dependent, with the latent variable $Z$ tied to specific network architectures and derived from intermediate activations, making them less flexible and non-stationary.
@@
PRG leverages the infinite-layer structure of continuous-time stochastic flow models [Chen2018], providing high expressiveness with relatively small parameter sizes. This structure underpins the success of modern generative models like SD3 [rombach2022high] and DALL·E 3 [betker2023improving]. By using a reversible process, our method allows for the simultaneous training of discriminative and generative models, offering excellent performance with enhanced capacity.
@@
Features extracted by reversing the generative process remain stable along the reverse trajectory, allowing feature use from any point $t\in[0,1]$ for downstream tasks, not just $x_{1}$. These features are robust to various discretization schemes (e.g., Euler, Runge-Kutta) and time steps ($10^{3}$ to $10^{1}$). Our method generalizes well across datasets, enabling fine-tuning of pretrained models for new tasks. Notably, it applies to all diffusion and flow models and has been successfully transferred to community-developed text-to-image models in LABEL:sec:trans.
A more detailed analysis of why PRG is effective can be found in Appendix A.
@@ 4.1 Settings
For generative model selection, we adopt three diffusion and flow-based models: GVP, ICFM, and OTCFM. For pre-training, we followed the protocol from [nichol2021improved], using the Adam optimizer [kingma2014adam] with a fixed learning rate of $1\times 10^{-4}$ for 1,200 epochs on the respective training sets. For downstream image classification, we adopted the configuration from [liu2021swin], using AdamW for 200 epochs with a cosine decay learning rate schedule and a 5-epoch linear warm-up. The training used a batch size of 128 and an initial learning rate of 0.001. We followed DDAE [xiang2023denoising]’s setting (64 resolution, simple data augmentation, no mixup [zhang2017mixup], no cutmix [yun2019cutmix]) for fair comparison, yet our method achieves SOTA diffusion classifier performance. Further details, including hyperparameters and efficiency statistics for training and inference, are provided in Secs. B.1, B.2 and B.3 .
@@
We select state-of-the-art (SOTA) representation models as baselines to evaluate PRG on CIFAR-10 [Krizhevsky2009LearningML], Tiny-ImageNet [Le2015TinyIV], and ImageNet [Russakovsky2014ImageNetLS]. Baselines include classical discriminative models like WideResNet-$28$-$10$ [zagoruyko2016wide] and ResNeXt-$29$-$16\times 64$[xie2017aggregated], as well as diffusion-based generative models such as GLOW [fetaya2019understanding], Energy Model [duvenaud2020your], SBGC [zimmermann2021score], HybViT [yang2022your], and DDAE [xiang2023denoising]. For out-of-distribution (OOD) evaluation, we compare PRG with PGD [madry2017towards], PLAT [wang2020enresnet], AugMix [hendrycks2019augmix], AutoAug [cubuk2019autoaugment], SBGC, and PDE+ [yuan2024pde] on CIFAR-10-C and Tiny-ImageNet-C. All results are from their original papers.